Circular No. 6879

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1998 G1 (LINEAR) Continuing orbit computations, mainly by S. Nakano and the under-

signed, show that this object is of Halley type. The orbital elements on MPEC 1998-H04 have P = 41.5 years. D. DeGraff remarks that his coaddition of CCD frames obtained on Apr. 6.1 UT by A. C. Schwortz and himself with the 0.81-m reflector at the Alfred University Observatory appears to confirm the presence of the tail reported on IAUC 6863. The feature was in p.a. ∼ 100^{o}, the shorter length of 10′′ probably being due to the presence of the moon nearby.

COMET C/1998 G2 (SOHO) C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discovery by D. Biesecker of a comet, tailless and no brighter than mag 6, in C2 coronagraphic data. This comet appears to have been obscured by the occulting disk pylon in the C3 coronagraph and was not detected. Details are on MPEC 1998-H05.

1998 UT  α_{2000}  δ_{2000}
Apr. 2.080  0^{h}52.^{m}2  + 4^{o}02′
1997 XF_{11}

The estimate by the undersigned on IAUC 6837 that passage within 0.002 AU of the earth on 2028 Oct. 26 was “virtually certain” was incorrect; this was a 1σ miss distance, and detailed computations allowed miss distances of up to 0.02–0.03 AU. The nominal miss distance and error by D. Yeomans and P. Chodas (see also IAUC 6839) should have been given as 0.00058_{−0.00039}^{+0.00892} AU (3σ on the plus side), and K. Muinonen early remarked on a 10-percent chance for a miss by more than the lunar distance. All concerned, including E. Bowell, agree that in the absence of effects that would be highly unusual there was no possibility of an earth encounter within 0.00019–0.00021 AU; Chodas remarks (and the undersigned agrees) that this was already evident with the issuance of MPEC 1997-Y11. Nevertheless, the likelihood of an unusually close approach in 2028 was also clear already in Dec. 1997, and it is unfortunate that there was not then greater awareness of this, for the recognition of 1990 observations (also found by Bowell on films taken by C. S. and E. M. Shoemaker) would have been possible at an earlier date and physical observations may have been attempted. Given the Helin-Lawrence 1990 measurements, simultaneous computations by G. V. Williams and by Chodas immediately made it clear to all concerned that the 2028 miss distance would be in the range 0.006–0.007 AU.

1998 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden